3. Calculate which of the following contains more grams of calcium: 4.6 X 1020 atoms of Calcium or 0.050 mol calcium.
1
Expert's answer
2021-10-30T11:16:17-0400

Molar mass of calcium = 40.078

= 0.05×40.078= 2.0039g

4.6×1020/6.023×1023=0.000764mol

0.000764×40.078=0.03116g

Hence, 0.050mol has more grams